A committee of 8 is to be selected out of 6 males and 8 females. In how many ways can it beformed so that the male way not be out numbered?
Answer by stanbon(75887)

A committee of 8 is to be selected out of 6 males and 8 females. In how many ways can it beformed so that the male way not be out numbered?
----
6males/2females: 6C6*8C2 = 1*28 = 28 ways
---
5males/3females: 6C5*8C3 = 6*56 = 336 ways
-----
4males/4females: 6C4*8C4 = 15*70 = 1050 ways
=====
Total # of ways: 28+336+1050 = 1414
========================================
Cheers,
Stan H.
===========